Three identical metal spheres are hung from silk threads so that they are isolated from their surroundings. Initially, they are neutral in charge. A positively charged rod is touched to sphere A, sphere A is then touched to sphere B, and sphere B is touched to sphere C. If the charge on the rod was initially 20 C, .



Which of the three spheres has the most protons?

If all three spheres are briefly and simultaneously touched together, then what will be the resulting charge on sphere B?

What is the charge of sphere B after it has been touched to sphere C?


Which of three spheres has the least electrons after sphere B touched to sphere C?


Total charge that is distributed among the 3 spheres?


Charge on sphere A after it is touched to sphere B?

Smallest charge that any of the 3 spheres can have after all of the spheres are carrying a charge?


Charge on sphere C after it has been touched by sphere B?


Charge on sphere A after it is touched by the rod?
1
Expert's answer
2020-02-20T09:58:25-0500

Since the spheres are identical, they have equal number of protons.


The total charge of the spheres is 20/2=10C Hence the resulting charge on sphere B is 10/3=3.33C


After A-B touch the charge of B is 10/2=5C Hence after B-C touch the charge of B is 5/2=2.5C


Sphere A has thee least electrons because it has the biggest positive charge.


Total charge distribured is 20/2=10C


After A-B touch the charge of A is 10/2=5C


2.5C is the smallest charge in this system because any following combination of touches will result in no charge decrease.


After B-C touch the charge of C is 5/2=2.5C


After A-rod touch the charge of A is 210/2=10C